# Muon beams towards muonium physics: progress and prospects

### v1

@@ III.2.1 Ionization cooling
As plotted in Fig. 13, the principle of this method is to reduce the muon energy through ionization losses in a low-Z absorber (such as liquid hydrogen or lithium hydride), then the muons are reaccelerated longitudinally in an RF cavity. By repeating this process, the transverse emittance of the muon beam can be compressed. The Muon Ionization Cooling Experiment (MICE), conducted at the ISIS muon source in the UK, reported the first experimental demonstration of ionization cooling [MICE:2019jkl].
@@ III.2.2 Friction cooling
Muons slowed though cold moderator still exhibit a very large phase space volume. PSI therefore proposed a friction cooling technique, as shown in Fig. 14 [Antognini:2021fae]. The muon beam enters a target region filled with cryogenic He gas, where strong electric and magnetic fields are applied. The muon energy rapidly decreases from the MeV scale to the eV scale through collisions with He. Once the energy reduces to the eV range, the electric field becomes dominant. Together with the magnetic field and density gradient of the gas, it induces a drift motion that compresses the broad spatial distribution and focuses the muon beam into a sub-mm size. The beam is then compressed along the -direction at room temperature. Finally, the muons are extracted from the gas into vacuum and re-accelerated along the direction, allowing the beam to exit the magnetic field with improved phase space quality. However, multiple scattering with the gas constrains the overall cooling performance. In 2020, the muCool collaboration demonstrated transverse compression, increasing the phase space density by a factor of with an efficiency of [Antognini:2020uyp].
@@ III.2.3 Muonium laser ionization cooling
Injecting surface muons into a tungsten foil heated to high temperature allows muonium to escape into vacuum with a kinetic energy of about 0.2 eV and an efficiency of 1% [Nagamine:1995zz]. A pulsed laser with an appropriate wavelength first excites the muonium atoms, and a second laser ionizes the excited state to produce slow muons. The principle of muonium laser ionization cooling method is shown as Fig. 15. The ionization efficiency can reach 100% in principle, although the polarization of the resulting slow muons is reduced to 50%. Note that only can be slowed by this method, as would be captured by nuclei.
In 1995, K. Nagamine et al. achieved the first slow-muon production using laser ionization of muonium, with a yield of 0.07 s [Nagamine:1995zz]. In 2008, P. Bakule et al. implemented the same technique at the ISIS muon source, increasing the yield to 15 s with an efficiency of the order of [Bakule:2008zz]. Compared to the LEM beamline at PSI, the USM beamline requires a pulsed muon source and operates with lower yield and polarization, but it provides superior beam quality and achieves lower energies. Fortunately, as introduced by Section IV.1, novel materials have been shown to substantially enhance both the production and emission rates of muonium, such as the currently adopted silica aerogel [Beare:2020gzr, Beer:2014ooa, Zhao:2023plv]. At present, the Ultra-Slow Muon (USM) beamline at the MUSE muon facility in Japan delivers a yield of 300 s [Kanda:2023gqp]. Future progress can be expected through further developments in both the laser system and the muonium formation materials.
@@ III.2.4 Muon acceleration
In 2018, S. Bae et al. conducted the first demonstration in muon acceleration at J-PARC MUSE [Bae:2018atj]. Negative muonium ions (, ) were produced via electron capture as surface muons passed through an aluminum degrader, and were accelerated from 50 keV to 89 keV using a radio-frequency quadrupole linac. Although the experiment served only as a proof of principle and was limited by the large emittance and extremely low yield of , it provided the first experimental evidence that RF structures can accelerate muons. Furthermore, a scheme of muon linac (up to 212 MeV) based on above techniques as shown as Fig. 16 is also proposed at the H line of MUSE [Kondo:2022sqb]. In 2025, S. Aritome et al. achieved a technical milestone by demonstrating RF acceleration of slow for the first time at J-PARC MUSE [Aritome:2024rlu]. Muonium laser ionization cooling was employed, followed by collection and focusing with an electrostatic lens system. The muons were then pre-accelerated to 5.7 keV by electrostatic fields and injected into an RFQ, where they were further accelerated to a final energy of 100 keV. The experiment showed that the transverse emittance was reduced by two orders of magnitude relative to the initial surface muon beam.
@@ IV.1 Muonium production
Muonium is first observed in a high-pressure argon gas environment by V.W. Hughes et al. in 1960 [Hughes:1960zz]. Because the muon beams at that time had relatively high momenta, the gas had to be compressed to 40 atmospheres to effectively stop the muons [Walker_1983]. The muoniums were observed using SR methods. The muonium formation efficiency in the rare gas can approach nearly 100%. However, since the electron spin distributes randomly, the muonium polarization is limited to 50% [Jungmann:2004sa]. Applying a strong magnetic field can increase the polarization to 100%. Experimental studies of muonium formation in various gases were carried out [Hughes:1970zza]. Later, muonium production in solids [Gurevich:1971ai] and liquids [Percival_1976] was also observed. The drawback of this method is that the muonium can hardly emit outside the target. Even in solid rare gas targets, only a very small fraction () of M can escape into vacuum [Prokscha:2008zz]. For precision measurements such as hyperfine structure splitting or the 1S-2S transition, it is necessary to eliminate the perturbations to the energy levels induced by collisions with the gas, thereby motivating the search for materials capable of emitting muonium into vacuum.
@@
With the invention of surface muon beams [Pifer:1976ia], studies of muonium production advanced significantly. Inspired by studies of positronium, several solid targets began to be employed for muonium formation [Canter_1974]. In 1985, A.P. Mills et al. observed muonium emission from 99.99% pure tungsten foil heated to high temperatures. The investigated temperatures ranged from 300 K to 2800 K, achieving a muonium yield in vacuum per muon of 4% at 2300 K [Mills:1986zzb]. However, this production method was later found to depend strongly on the purity of the tungsten [Matsushita_1996]. Notably, PSI have developed a muonium source based on the slow muons from E4 beamline recently. To generate a sufficient number of metastable muonium for improving the precision of Lamb shift and fine structure measurements, the beam-foil method is required [Andr__1974]. In earlier experiments at TRIUMF and LAMPF, this approach needed substantial muon degradation [C_J_Oram_1981, Badertscher_1992], which reduced both the M(2S) yield and the overall experimental precision. In 2020, G. Janka et al. employed a 15 nm carbon foil to stop slow muons provided by the LEM beamline, achieving a muonium (1S and 2S) yield in vacuum of 31.8% for 10 keV incident muons, where M(2S) accounted for 10% of the total muonium yield [Janka:2020xky]. This technique is expected to improve the precision of Lamb shift measurements by two orders of magnitude.
@@
Numerous experiments explored the use of various silica-based solid targets for muonium production. In 1992, W. Schwarz et al. evaluated the performance of silica powder and aerogel targets at PSI and RAL [Schwarz_1992]. The experimental data were globally fit to a diffusion model to obtain the total yield in vacuum. Silica powder showed a clear advantage: the powder target achieved 8.2%, whereas the optimized aerogel target got 2.3%. Previous measurements indicated a muonium formation efficiency of 60% in silica powder, reported by R.F. Kiefl in 1979 [Kiefl_1979], while W. Schwarz et al. reported formation efficiency of 40–50% for aerogels of varying densities. However, when considering overall performance, silica aerogel proved to be the superior option. Its self-supporting structure eliminates the scattering introduced by the supporting foil required for powder targets, enabling the production of higher-quality beams for precision applications such as laser-ionized ultra-slow muon sources introduced in Section III.2.3. This work not only established the advantages of aerogel as a stable and self-supporting target material but also demonstrated that specialized surface treatments can further enhance its yield.
@@
However, for porous materials above, muonium reaches the vacuum via thermal diffusion, causing broad energy and angular distributions. Taking advantage of its unique properties, superfluid helium (SFHe) has demonstrated potential as a next-generation muonium formation target. When impurity atoms or ions enter SFHe, they are repelled by the Pauli repulsion and form a surrounding cavity inside the material as plotted in Fig. 17; once reaching the surface, they will be expelled by the chemical potential to achieve an almost monoenergetic emission of meV [Saarela_1993, Hayden_1995, Mar_n_1998]. Because the ionization energy of He atoms is far higher than the binding energy of muonium, muonium formation is not observed in either gaseous or liquid He. In 1992, E. Krasnoperov et al. conducted experiments with a SR spectrometer on the M3 beamline at PSI. The precession signal of triplet-state muonium in SFHe below 1 K was observed, thereby demonstrating that SFHe can serve as a medium for muonium formation for the first time [Krasnoperov_1992].
@@ IV.2 Muonium: charged lepton flavor violation
Taking advantage of the rapid development of particle detection techniques and the construction of muon beams in China, Jian Tang et al. have proposed a new experiment aiming to search for the conversion with a higher sensitivity, known as the Muonium-to-Antimuonium Conversion Experiment (MACE) [Bai:2022sxq, Zhao:2024qjb, Bai:2024skk, An:2025lws]. The conceptual design of the detector system is as described in Fig. 18. MACE will utilize a surface muon beam with a momentum of 24 MeV/ and an on-target intensity of /s. The detector system consists of a Michel electron magnetic spectrometer (MMS), a momentum-selection positron transport system (PTS), and a positron detection system (PDS). In the conceptual design, subdetectors with higher precision were designed and validated by full simulation [Zhao:2024qjb, Bai:2024skk, Chen:2024jmg, Lu:2025col]. Particularly, a silica aerogel target is proposed to increase the muonium yield rate by a factor of 2, highly improving the statistical precision [Zhao:2023plv]. MACE aims to constrain the conversion probability to the level, which is two orders of magnitude better than the previous MACS at PSI. Further improvement on the sensitivity is expected to be achieved by the enhancement of muonium yield of the target and the resolution of each subdetector system.
@@ IV.3 Precise muonium spectroscopy
where the is the reduced mass of the muonium system, is the Coulomb potential. According to the Eq. 8, the ground state wave function and the ionization energy of muonium can be expressed as
@@
where is the hyperfine coupling constant. As a reference, Fig. 1 shows the muonium energy levels, including the hyperfine interaction. For the free muonium, the polarization is decoupled from time evolution. Because the accelerator muon beam is nearly polarized, and the electron spin in the material is almost isotropic, the muonium states and are produced with equal probability.
@@
Under the external magnetic field, the total spin eigenstates of muonium are no longer good. Instead, the eigenstates of muonium can be expressed as a linear combination of Zeeman states,
@@ IV.3.1 Muonium Lamb shift
Both TRIUMF and LAMPF conducted measurements of the Lamb shift in muonium, reporting results of MHz [Oram:1983sd] and MHz [Woodle:1990ky], respectively. The precision of these measurements was primarily limited by low statistics and high background levels. In 2021, the MuoniuM lAser SpectroScopy (Mu-MASS) experiment as shown as Fig. 20 at PSI reported the most recent result of MHz [Mu-MASS:2021uou], representing an order of magnitude improvement comparing to the previous measurement. The experiment utilized surface muons from the E4 beamline. Muons were first slowed down to 20 eV by a Ne moderator, then re-accelerated to 7.5 kV and formed muonium atoms in a 10 nm thick carbon foil target. Part of the muonium atoms were excited to the long-lived 2S state. A static electric field of 250 V/cm mixed the 2S and 2P states, relaxing them to the ground state within a few ns, accompanied by the emission of 122 nm Lyman-alpha photons. When microwave radiation resonant with the transition was applied, it reduced the population of 2S muonium atoms, thereby reducing the Lyman-alpha photon signal. By scanning the microwave frequency around the transition, the Lamb shift can be determined by the centroid energy difference between and averaged over hyperfine structure.
@@ IV.3.2 Muonium 1S-2S transition
However, the chirping effect of pulsed laser beams still significantly limits the precision of measurement. With development in laser technology, CW laser beams have become available to reduce the chirping effect, leading to a possible improvement of two orders of magnitude. Furthermore, if Doppler shifts and AC Stark shifts can be corrected, the precision could be improved by an additional order of magnitude [Ohayon:2021dec]. Ref. [Zhadnov:2023vcr] provides a detailed introduction of the laser system setup. The Mu-MASS experiment at PSI will produce muonium in a porous silica film target at 100 K using low-energy muons from the E4 beamline. The experimental setup can be seen in Fig. 21. Under these conditions, the muonium formation efficiency can reach approximately 20%. Taking advantage of the advanced laser techniques and a new detection scheme from the positronium 1S-2S spectroscopy, the uncertainty in the transition probability measurement is expected to be reduced to the level of 10 kHz [Crivelli:2018vfe, Cortinovis:2023zqi].
@@ IV.3.3 Muonium hyperfine structure
In 1999, W. Liu et al. reported the most precise measurement of the muonium hyperfine structure interval under the high-field condition [Liu:1999iz]. The experiment was carried out at the LAMPF. Muons were directed into a strong magnetic field of about 1.7 T. At the center, a copper microwave resonant cavity was filled with high-purity krypton gas (pressure of 0.8 or 1.5 atm) as the muonium production target. In the strong magnetic field, the ground-state energy levels of muonium splits caused by Zeeman effect. A transverse microwave field was applied, and its frequency was finely tuned to match Zeeman transitions and . The muon spin flipped, causing the angular distribution asymmetry of the Michel positrons. The positrons were detected by a scintillator counter downstream of the cavity, and the resonance signal was constructed from the difference in counts with the microwave on and off. In the traditional method, the resonance linewidth is limited by the muon lifetime ( kHz). This experiment applied a resonance line narrowing technique [Liu:1995hq]. The muon beam was time-modulated by a chopper (beam period with 4 s on, 10 s off), and Michel positrons were detected in eleven time windows of 0.95 s. The “old” muonium survived several times longer than the muon lifetime. The produced linewidth is three times narrower and stronger in amplitude than in the previous experiments. This innovation greatly improved the precision of measurement, resulting in the muonium hyperfine structure MHz (12 ppb), the magnetic moment ratio (120 ppb) [Liu:1999iz]. These measured value could deduce the (27 ppb), and the fine structure constant (58 ppb) [Liu:1999iz].
@@
In 2021, the MuSEUM (Muonium Spectroscopy Experiment Using Microwave) collaboration reported a measurement of the muonium hyperfine structure [Kanda:2020mmc]. The experiment was conducted at the high-intensity pulsed muon source of J-PARC, as shown in Fig. 22. Muons were stopped and formed muonium in a krypton gas chamber at 1 bar. The hyperfine structure corresponds to the energy splitting between the spin-singlet state (25%) and spin-triplet state (75%) of muonium. An oxygen-free copper resonant cavity generated a microwave field, which could be tuned in the range of 4461–4464.8 MHz by inserting a movable aluminum rod in the chamber. Magnetic shielding was applied to keep a near-zero field. Downstream of the chamber, a two-layer plastic scintillator array coupled to SiPMs was used for coincidence measurements of Michel positrons. The final measured value of the muonium HFS is 4463.302(4) MHz, with a relative precision of 0.9 ppm [Kanda:2020mmc]. A subsequent study introduced a new method, Rabi-oscillation spectroscopy, which improves the precision by a factor of two compared with conventional frequency-swept spectroscopy. By taking advantage of this more efficient measurement, the result reaches 160 ppb, which is the highest precision under zero magnetic field to date [MuSEUM:2020mzm]. In the future, the MuSEUM experiment aims to achieve 12 ppb under zero-field conditions and 1.2 ppb under high-field conditions with a higher intensity muon beam [MuSEUM:2025cmo].

### v2

@@ III.2.1 Ionization cooling
As plotted in Fig.˜13, the principle of this method is to reduce the muon energy through ionization losses in a low-Z absorber (such as liquid hydrogen or lithium hydride), then the muons are reaccelerated longitudinally in an RF cavity. By repeating this process, the transverse emittance of the muon beam can be compressed. The Muon Ionization Cooling Experiment (MICE), conducted at the ISIS muon source in the UK, reported the first experimental demonstration of ionization cooling [MICE:2019jkl].
@@ III.2.2 Friction cooling
Muons slowed through a cold moderator still exhibit a very large phase space volume. PSI therefore proposed a friction cooling technique, as shown in Fig.˜14 [Antognini:2021fae]. The muon beam enters a target region filled with cryogenic He gas, where strong electric and magnetic fields are applied. The muon energy rapidly decreases from the MeV scale to the eV scale through collisions with He. Once the energy reduces to the eV range, the electric field becomes dominant. Together with the magnetic field and density gradient of the gas, it induces a drift motion that compresses the broad spatial distribution and focuses the muon beam into a sub-mm size. The beam is then compressed along the -direction at room temperature. Finally, the muons are extracted from the gas into vacuum and re-accelerated along the direction, allowing the beam to exit the magnetic field with improved phase space quality. However, multiple scattering with the gas constrains the overall cooling performance. In 2020, the muCool collaboration demonstrated transverse compression, increasing the phase space density by a factor of with an efficiency of [Antognini:2020uyp].
@@ III.2.3 Muonium laser ionization cooling
Injecting surface muons into a tungsten foil heated to high temperature allows muonium to escape into vacuum with a kinetic energy of about 0.2 eV and an efficiency of 1% [Nagamine:1995zz]. A pulsed laser with an appropriate wavelength first excites the muonium atoms, and a second laser ionizes the excited state to produce slow muons. The principle of muonium laser ionization cooling method is shown as Fig.˜15. The ionization efficiency can reach 100% in principle, although the polarization of the resulting slow muons is reduced to 50%. Note that only can be slowed by this method, as would be captured by nuclei.
In 1995, K. Nagamine et al. achieved the first slow-muon production using laser ionization of muonium, with a yield of 0.07 s [Nagamine:1995zz]. In 2008, P. Bakule et al. implemented the same technique at the ISIS muon source, increasing the yield to 15 s with an efficiency of the order of [Bakule:2008zz]. Compared to the LEM beamline at PSI, the USM beamline requires a pulsed muon source and operates with lower yield and polarization, but it provides superior beam quality and achieves lower energies. Fortunately, as introduced by Section˜IV.1, novel materials have been shown to substantially enhance both the production and emission rates of muonium, such as the currently adopted silica aerogel [Beare:2020gzr, Beer:2014ooa, Zhao:2023plv]. At present, the Ultra-Slow Muon (USM) beamline at the MUSE muon facility in Japan delivers a yield of 300 s [Kanda:2023gqp]. Future progress can be expected through further developments in both the laser system and the muonium formation materials.
@@ III.2.4 Muon acceleration
In 2018, S. Bae et al. conducted the first demonstration of muon acceleration at J-PARC MUSE [Bae:2018atj]. Negative muonium ions (, ) were produced via electron capture as surface muons passed through an aluminum degrader, and were accelerated from 50 keV to 89 keV using a radio-frequency quadrupole linac. Although the experiment served only as a proof of principle and was limited by the large emittance and extremely low yield of , it provided the first experimental evidence that RF structures can accelerate muons. Furthermore, a scheme of a muon linac (up to 212 MeV) based on above techniques as shown in Fig.˜16 is also proposed at the H line of MUSE [Kondo:2022sqb]. In 2025, S. Aritome et al. achieved a technical milestone by demonstrating RF acceleration of slow for the first time at J-PARC MUSE [Aritome:2024rlu]. Muonium laser ionization cooling was employed, followed by collection and focusing with an electrostatic lens system. The muons were then pre-accelerated to 5.7 keV by electrostatic fields and injected into an RFQ, where they were further accelerated to a final energy of 100 keV. The experiment showed that the transverse emittance was reduced by two orders of magnitude relative to the initial surface muon beam.
@@ IV.1 Muonium production
Muonium was first observed in a high-pressure argon gas environment by V.W. Hughes et al. in 1960 [Hughes:1960zz]. Because the muon beams at that time had relatively high momenta, the gas had to be compressed to 40 atmospheres to effectively stop the muons [Walker_1983]. The muoniums were observed using SR methods. The muonium formation efficiency in the rare gas can approach nearly 100%. However, since the electron spin distributes randomly, the muonium polarization is limited to 50% [Jungmann:2004sa]. Applying a strong magnetic field can increase the polarization to 100%. Experimental studies of muonium formation in various gases were carried out [Hughes:1970zza]. Later, muonium production in solids [Gurevich:1971ai] and liquids [Percival_1976] was also observed. The drawback of this method is that the muonium can hardly emit outside the target. Even in solid rare gas targets, only a very small fraction () of M can escape into vacuum [Prokscha:2008zz]. For precision measurements such as hyperfine structure splitting or the 1S-2S transition, it is necessary to eliminate the perturbations to the energy levels induced by collisions with the gas, thereby motivating the search for materials capable of emitting muonium into vacuum.
@@
With the invention of surface muon beams [Pifer:1976ia], studies of muonium production advanced significantly. Inspired by studies of positronium, several solid targets began to be employed for muonium formation [Canter_1974]. In 1985, A.P. Mills et al. observed muonium emission from 99.99% pure tungsten foil heated to high temperatures. The investigated temperatures ranged from 300 K to 2800 K, achieving a muonium yield in vacuum per muon of 4% at 2300 K [Mills:1986zzb]. However, this production method was later found to depend strongly on the purity of the tungsten [Matsushita_1996]. Notably, PSI has developed a muonium source based on the slow muons from E4 beamline recently. To generate a sufficient number of metastable muonium for improving the precision of Lamb shift and fine structure measurements, the beam-foil method is required [Andr__1974]. In earlier experiments at TRIUMF and LAMPF, this approach required substantial muon degradation [C_J_Oram_1981, Badertscher_1992], which reduced both the M(2S) yield and the overall experimental precision. In 2020, G. Janka et al. employed a 15 nm carbon foil to stop slow muons provided by the LEM beamline, achieving a muonium (1S and 2S) yield in vacuum of 31.8% for 10 keV incident muons, where M(2S) accounted for 10% of the total muonium yield [Janka:2020xky]. This technique is expected to improve the precision of Lamb shift measurements by two orders of magnitude.
@@
Numerous experiments explored the use of various silica-based solid targets for muonium production. In 1992, W. Schwarz et al. evaluated the performance of silica powder and aerogel targets at PSI and RAL [Schwarz_1992]. The experimental data were globally fit to a diffusion model to obtain the total yield in vacuum. Silica powder showed a clear advantage: the powder target achieved 8.2%, whereas the optimized aerogel target got 2.3%. Previous measurements indicated a muonium formation efficiency of 60% in silica powder, reported by R.F. Kiefl in 1979 [Kiefl_1979], while W. Schwarz et al. reported formation efficiency of 40–50% for aerogels of varying densities. However, when considering overall performance, silica aerogel proved to be the superior option. Its self-supporting structure eliminates the scattering introduced by the supporting foil required for powder targets, enabling the production of higher-quality beams for precision applications such as laser-ionized ultra-slow muon sources introduced in Section˜III.2.3. This work not only established the advantages of aerogel as a stable and self-supporting target material but also demonstrated that specialized surface treatments can further enhance its yield.
@@
However, for porous materials above, muonium reaches the vacuum via thermal diffusion, causing broad energy and angular distributions. Taking advantage of its unique properties, superfluid helium (SFHe) has demonstrated potential as a next-generation muonium formation target. When impurity atoms or ions enter SFHe, they are repelled by the Pauli repulsion and form a surrounding cavity inside the material as plotted in Fig.˜17; once reaching the surface, they will be expelled by the chemical potential to achieve an almost monoenergetic emission of meV [Saarela_1993, Hayden_1995, Mar_n_1998]. Because the ionization energy of He atoms is far higher than the binding energy of muonium, muonium formation is not observed in either gaseous or liquid He. In 1992, E. Krasnoperov et al. conducted experiments with a SR spectrometer on the M3 beamline at PSI. The precession signal of triplet-state muonium in SFHe below 1 K was observed, thereby demonstrating that SFHe can serve as a medium for muonium formation for the first time [Krasnoperov_1992].
@@ IV.2 Muonium: charged lepton flavor violation
Taking advantage of the rapid development of particle detection techniques and the construction of muon beams in China, Jian Tang et al. have proposed a new experiment aiming to search for the conversion with a higher sensitivity, known as the Muonium-to-Antimuonium Conversion Experiment (MACE) [Bai:2022sxq, Zhao:2024qjb, Bai:2024skk, An:2025lws]. The conceptual design of the detector system is as described in Fig.˜18. MACE will utilize a surface muon beam with a momentum of 24 MeV/ and an on-target intensity of /s. The detector system consists of a Michel electron magnetic spectrometer (MMS), a momentum-selection positron transport system (PTS), and a positron detection system (PDS). In the conceptual design, subdetectors with higher precision were designed and validated by full simulation [Zhao:2024qjb, Bai:2024skk, Chen:2024jmg, Lu:2025col]. Particularly, a silica aerogel target is proposed to increase the muonium yield rate by a factor of 2, highly improving the statistical precision [Zhao:2023plv]. MACE aims to constrain the conversion probability to the level, which is two orders of magnitude better than the previous MACS at PSI. Fig.˜19 illustrates the new-physics sensitivity achieved by past experiments as well as that anticipated for the MACE project. Further improvement on the sensitivity is expected to be achieved by the enhancement of muonium yield of the target and the resolution of each subdetector system.
@@ IV.3 Precise muonium spectroscopy
where is the reduced mass of the muonium system, is the Coulomb potential. According to the Eq.˜8, the ground state wave function and the ionization energy of muonium can be expressed as
@@
where is the hyperfine coupling constant. As a reference, Fig.˜1 shows the muonium energy levels, including the hyperfine interaction. For the free muonium, the polarization is decoupled from time evolution. Because the accelerator muon beam is nearly polarized, and the electron spin in the material is almost isotropic, the muonium states and are produced with equal probability.
@@
Under the external magnetic field, the eigenstates of muonium can be expressed as a linear combination of Zeeman states,
@@ IV.3.1 Muonium Lamb shift
Both TRIUMF and LAMPF conducted measurements of the Lamb shift in muonium, reporting results of MHz [Oram:1983sd] and MHz [Woodle:1990ky], respectively. The precision of these measurements was primarily limited by low statistics and high background levels. In 2021, the MuoniuM lAser SpectroScopy (Mu-MASS) experiment as shown in Fig.˜20 at PSI reported the most recent result of MHz [Mu-MASS:2021uou], representing an order of magnitude improvement comparing to the previous measurement. The experiment utilized surface muons from the E4 beamline. Muons were first slowed down to 20 eV by a Ne moderator, then re-accelerated to 7.5 kV and formed muonium atoms in a 10 nm thick carbon foil target. Part of the muonium atoms were excited to the long-lived 2S state. A static electric field of 250 V/cm mixed the 2S and 2P states, relaxing them to the ground state within a few ns, accompanied by the emission of 122 nm Lyman-alpha photons. When microwave radiation resonant with the transition was applied, it reduced the population of 2S muonium atoms, thereby reducing the Lyman-alpha photon signal. By scanning the microwave frequency around the transition, the Lamb shift can be determined by the centroid energy difference between and averaged over hyperfine structure.
@@ IV.3.2 Muonium 1S-2S transition
However, the chirping effect of pulsed laser beams still significantly limits the precision of measurement. With development in laser technology, CW laser beams have become available to reduce the chirping effect, leading to a possible improvement of two orders of magnitude. Furthermore, if Doppler shifts and AC Stark shifts can be corrected, the precision could be improved by an additional order of magnitude [Ohayon:2021dec]. Ref. [Zhadnov:2023vcr] provides a detailed introduction of the laser system setup. The Mu-MASS experiment at PSI will produce muonium in a porous silica film target at 100 K using low-energy muons from the E4 beamline. The experimental setup can be seen in Fig.˜21. Under these conditions, the muonium formation efficiency can reach approximately 20%. Taking advantage of the advanced laser techniques and a new detection scheme from the positronium 1S-2S spectroscopy, the uncertainty in the transition probability measurement is expected to be reduced to the level of 10 kHz [Crivelli:2018vfe, Cortinovis:2023zqi]. Additionally, it’s worth mentioning that the Advanced Muonium Laser Experiment at Tokai (AmuLET) collaboration has preliminarily reported the first measurement of the transition recently [UetakeSSP2025].
@@ IV.3.3 Muonium hyperfine structure
In 1999, W. Liu et al. reported the most precise measurement of the muonium hyperfine structure interval under the high-field condition [Liu:1999iz]. The experiment was carried out at the LAMPF. Muons were directed into a strong magnetic field of about 1.7 T. At the center, a copper microwave resonant cavity was filled with high-purity krypton gas (pressure of 0.8 or 1.5 atm) as the muonium production target. In the strong magnetic field, the ground-state energy levels of muonium splits caused by Zeeman effect. A transverse microwave field was applied, and its frequency was finely tuned to match Zeeman transitions and . The muon spin flips, causing the angular distribution asymmetry of the Michel positrons. The positrons were detected by a scintillator counter downstream of the cavity, and the resonance signal was constructed from the difference in counts with the microwave on and off. In the traditional method, the resonance linewidth is limited by the muon lifetime ( kHz). This experiment applied a resonance line narrowing technique [Liu:1995hq]. The muon beam was time-modulated by a chopper (beam period with 4 s on, 10 s off), and Michel positrons were detected in eleven time windows of 0.95 s. The “old" muonium survived several times longer than the muon lifetime. The produced linewidth is three times narrower and stronger in amplitude than in the previous experiments. This innovation greatly improved the precision of measurement, resulting in the muonium hyperfine structure MHz (12 ppb), the magnetic moment ratio (120 ppb) [Liu:1999iz]. These measured values could deduce the (27 ppb), and the fine structure constant (58 ppb) [Liu:1999iz].
@@
In 2021, the MuSEUM (Muonium Spectroscopy Experiment Using Microwave) collaboration reported a measurement of the muonium hyperfine structure [Kanda:2020mmc]. The experiment was conducted at the high-intensity pulsed muon source of J-PARC, as shown in Fig.˜22. Muons were stopped and formed muonium in a krypton gas chamber at 1 bar. The hyperfine structure corresponds to the energy splitting between the spin-singlet state (25%) and spin-triplet state (75%) of muonium. An oxygen-free copper resonant cavity generated a microwave field, which could be tuned in the range of 4461–4464.8 MHz by inserting a movable aluminum rod in the chamber. Magnetic shielding was applied to keep a near-zero field. Downstream of the chamber, a two-layer plastic scintillator array coupled to SiPMs was used for coincidence measurements of Michel positrons. The final measured value of the muonium HFS is 4463.302(4) MHz, with a relative precision of 0.9 ppm [Kanda:2020mmc]. A subsequent study introduced a new method, Rabi-oscillation spectroscopy, which improves the precision by a factor of two compared with conventional frequency-swept spectroscopy. By taking advantage of this more efficient measurement, the result reaches 160 ppb, which is the highest precision under zero magnetic field to date [MuSEUM:2020mzm]. In the future, the MuSEUM experiment aims to achieve 12 ppb under zero-field conditions and 1.2 ppb under high-field conditions with a higher intensity muon beam [MuSEUM:2025cmo].